OpenStudy (hba):

Integrate:

OpenStudy (hba):

\[\int\limits_{}^{}(x+1)(x+3)dx\]

OpenStudy (anonymous):

\[\int\limits_{?}^{?}{x^2+4x+3} = \left[ {1 \over 3}x^3 +2x^2 +3x + C \right]\]

hartnn (hartnn):

you are starting with integration ?

OpenStudy (hba):

Just Revising